OpenStudy (anonymous):

A common inhabitant of human intestines is the bacterium Escherichia coli, named after the German pediatrician Theodor Escherich, who identified it in 1885. A cell of this bacterium in a nutrient-broth medium divides into two cells every 20 minutes. The initial population of a culture is 40 cells. (a) Find the relative growth rate in hours.

OpenStudy (anonymous):

What we are given initial # of cells = 40. doubles every 20 minutes.

OpenStudy (amistre64):

how many are there after 1 hour? after 2 hours? after 3 hours?

OpenStudy (amistre64):

or rather, after 3 periods, after 6 periods, after 9 periods .... 3 periods of 20min in an hour right?

OpenStudy (anonymous):

no it just wants us to find the general relative growth rate

OpenStudy (amistre64):

then youll have to define the terminology to see if my idea is on the right track or not

OpenStudy (anonymous):

check if I did my work correct, one sec

OpenStudy (anonymous):

\[y=40e ^{kt}\] \[80=40e ^{k \frac{ 20 }{ 60 }}\] \[2=e ^{k \frac{ 1 }{ 3 }}\] \[\ln 2=\ln e ^{k \frac{ 1 }{ 3 }}\] \[k=3\ln 2\]

OpenStudy (amistre64):

is this calculus work?

OpenStudy (anonymous):

according to my book k is the relative growth rate. we are using Calculus Early Transcedentals 8 edition. and yes

OpenStudy (amistre64):

\[\frac{dP}{dt}=kP\] \[k(\color{red}{rgr})=\frac{1}{P}\frac{dP}{dt}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So what does it mean?

OpenStudy (amistre64):

P(0min)= 40 P(20min)= 80 P(40min)= 160 P(60min) = 320 20'; 640 40'; 1280 60'; 2560 P(0 hours) = 40 P(1 hours) = 320 P(2 hours) = 2560 320 = 40 r , r = 8 2560 = 40 8^3 ?? ... yes P(t) = 40 8^t in terms of hours, as opposed to 20minutes

OpenStudy (anonymous):

the question is asking to solve the value of k

OpenStudy (amistre64):

yeah, i was trying to work out a similar problem, but this might be a better approach.

OpenStudy (amistre64):

some texts are easier to read thru than others fer sure :)

OpenStudy (anonymous):

ah ok I was correct in my procedure then

OpenStudy (amistre64):

your process does match with their method yes :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

oh also one mroe thing

OpenStudy (anonymous):

if we want to find the rate of growth, we take the derivative of y=40e^(tln8) right?

OpenStudy (amistre64):

P(t) = 40 8^t P'(t) =40 ln(8) 8^t gives us k=ln(8) = 3ln(2)

OpenStudy (amistre64):

that was the other one i was looking at .... finally clicked as to where to get k from :) but the site is confusing ... which is why i asked for definitions of the terms to start with, ive been far removed from the content.

OpenStudy (anonymous):

oh ok lol

OpenStudy (amistre64):

it seems that k is the growth rate, ... and the relative growth rate?

OpenStudy (anonymous):

no in the book it says k is the relative growth rate, I dont know what the difference is to be honest.

OpenStudy (anonymous):

oh wait that just gave me the answer lol, dp/dt is the growth rate

OpenStudy (amistre64):

then is sounds like dP/dt is the growth rate since it is the thing being divded by P

OpenStudy (anonymous):

I dont know why but when im working alone, I can't seem to think thats why I like coming on here. It helps me learn haha

OpenStudy (amistre64):

but isnt that just solving for k?

OpenStudy (amistre64):

pfft ... getting wires crossed lol

OpenStudy (anonymous):

no dP/dt is not solving for k, it is the rate at which the cells are doubling in terms of time

OpenStudy (amistre64):

dP/dt is being called the growth rate. but that is just another function: dP/dt = k P

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so if we want to find derivative of P(t) we just take the product of k and P

OpenStudy (amistre64):

P = 40 8^t based on t=hours dP/dt = ln(8) 40 8^t yes

OpenStudy (anonymous):

can you show the steps you took to get 8 as the base plz?

OpenStudy (amistre64):

already did ... its posted above

OpenStudy (anonymous):

oh yea ok

OpenStudy (anonymous):

Well thank you so much again

OpenStudy (amistre64):

youre welcome, hopefully it becomes more clear :) good luck

OpenStudy (ckaranja):

It increases by four times The function is h=4t